# Multilayer network science: theory, methods, and applications

### v1

@@ Abstract
Multilayer network science has emerged as a central framework for analysing interconnected and interdependent complex systems. Its relevance has grown substantially with the increasing availability of rich, heterogeneous data, which makes it possible to uncover and exploit the inherently multilayered organisation of many real-world networks. In this review, we summarise recent developments in the field. On the theoretical and methodological front, we outline core concepts and survey advances in community detection, dynamical processes, temporal networks, higher-order interactions, and machine-learning-based approaches. On the application side, we discuss progress across diverse domains, including interdependent infrastructures, spreading dynamics, computational social science, economic and financial systems, ecological and climate networks, science-of-science studies, network medicine, and network neuroscience. We conclude with a forward-looking perspective, emphasizing the need for standardized datasets and software, deeper integration of temporal and higher-order structures, and a transition toward genuinely predictive models of complex systems.
@@ I Introduction
During the last 25 years, network science has established itself as a powerful and seamlessly integrated research discipline dorogovtsev03b; vespignani04; caldarelli07; barrat08; cohen10; newman10; estrada11b; estrada15; barabasi16; latora17; menczer20; bagrow24. At variance with other, more niched disciplines, it has allowed us to enrich and to synergize with other branches of science: from sociology and economics to biology, chemistry, and physics. This has been possible because of the data revolution and increased computational power that took place in the twilight of the 20th century. Nowadays, applications of network science tools pervade all fields of science but, most importantly, the discipline has changed the way we think about interactions in natural and human-made complex systems.
@@
The momentum for the development of network science came with the discovery that seemingly diverse networks have universal properties that pervade many natural and artificial systems. Since 1998, ignited by the works of Watts and Strogatz and Barabási and Albert Watts1998Jun; Barabasi1999Oct, systems as diverse as electric power grids, food chains, brain networks, protein networks, transcriptional networks, and social networks have been cast into the mathematical language of networks. We are now familiar with jargon that includes centrality, homogeneous and heterogeneous systems, communities, structural and dynamic phase transitions, weak and strong links, etc.
@@
Yet, as the field matured, it became increasingly clear that most real systems cannot be faithfully represented as individual networks. Many natural, social, and technological systems are instead composed of several types of interactions, giving rise to multilayer or interdependent structures kivela14; boccaletti14; battiston14; bianconi18; Aleta2019Mar; artime22; dedomenico23. The conceptual roots of this idea trace back to the sociological notion of multiplexity, where individuals were known to maintain several types of social ties simultaneously within the same population Wasserman1994Nov. These early multi-relational studies anticipated many of the ideas later formalized in network science. Then, with the rise of large-scale digital data around 2010, empirical analyses began to demonstrate that distinct layers of interaction can exhibit markedly different structural and dynamical patterns, underscoring the need for explicit multilayer representations Szell2010Aug.
@@
A major turning point came with the study of interdependent networks, where nodes in one layer depend on nodes in another layer for functionality buldyrev_2010; Vespignani2010Apr. This framework revealed that coupling multiple networks can fundamentally alter systemic behaviour, producing abrupt percolation transitions and cascading failures, phenomena not observed in isolated networks. Following these foundational insights, the field rapidly matured from focusing primarily on cascades and robustness to developing a comprehensive mathematical formalism and exploring a wide array of dynamic processes. Researchers worked to create a unified language and a robust mathematical framework to describe the diverse architectures of multilayer systems kivela14; boccaletti14; battiston14, enabling the extension of classical concepts of network science to the multilayer context, such as centrality measures, among others. This theoretical consolidation sparked extensive research on dynamics across layers, including diffusion, synchronization, epidemic spreading, and evolutionary games, revealing that interlayer coupling can either enhance or inhibit dynamical processes depending on structural correlations.
@@
In the following years, multilayer methodologies spread rapidly across disciplines. As we will discuss in the following sections, applications now span social, biological, and technological systems. Parallel advances in community detection and inference have made it possible to extract meaningful mesoscale organization from high-dimensional, heterogeneous data. More recent developments integrate temporal, categorical, and hierarchical aspects within multilayer frameworks and connect them to higher-order network representations such as hypergraphs. Today, the multilayer paradigm stands as one of the most powerful generalizations in network science, bridging theoretical, computational, and empirical approaches to the study of interconnected complex systems.
@@ II.1.1 Definitions and Formalisms
A prominent example of a multilayer system is the coupling of a power grid and the internet. One layer represents the power grid, with nodes as power stations and substations, while links represent transmission lines, illustrating the distribution of electrical power. Meanwhile, the other layer represents the internet, with nodes as data centres, routing hubs, and servers, linked by data transmission pathways. This setup highlights the interdependency between layers: the internet needs power to function, while modern power grids rely on internet-based communication for their operation, something that can be encoded in the interlayer links. This multilayer network model is essential for studying cascading failures and the overall resilience of interconnected systems buldyrev_2010; DeDomenico_2023.
A multiplex network is a specific and widely used type of multilayer network where nodes represent the same entity on every layer, and interlayer links only connect a node to itself across different layers. Each layer then represents a different type of relationship among the nodes. For instance, a social network can be modelled as a multiplex network where layers represent different platforms or social contexts kivela14.
To analyse multiplex networks, we need to generalize the adjacency matrix. Two primary formalisms are used. First, we have the tensor formalism, in which the separation of layers can be maintained explicitly. A multiplex network can then be represented by an adjacency tensor $\mathcal{M}$ of order 4 with elements $M_{ij}^{\alpha\beta}$. Here, $M_{ij}^{\alpha\alpha}$ represents the weight of the intralayer link between nodes $i$ and $j$ in layer $\alpha$, while $M_{ij}^{\alpha\beta}$ with $\alpha\neq\beta$ represents the interlayer connections. This is a more general and flexible representation, though it requires methods from tensor algebra for analysis DeDomenico2013.
Second, we have the formalism based on the supra-adjacency matrix, which represents the entire multilayer system as a single, larger network. For a multiplex network with $N$ nodes and $L$ layers, the supra-adjacency matrix $\mathcal{A}$ is a $NL\times NL$ block matrix. The diagonal blocks are the adjacency matrices of each layer $\alpha$, $\text{{A}}^{\alpha}$, representing intralayer connections. The off-diagonal blocks encode the interlayer connections, which can be represented by the identity matrix I in the $\alpha\beta$ block if layer $\alpha$ and $\beta$ are connected, and 0 otherwise kivela14. An example of supra-adjacency matrix of a multiplex network is
@@
Following these representations, we can also define the corresponding Laplacian matrices. For a single-layer graph, the Laplacian matrix is defined as $\text{{L}}=\text{{D}}-\text{{A}}$, where D is the diagonal matrix of node degrees. Similarly, the supra-Laplacian matrix $\mathcal{L}$ can be constructed from the supra-adjacency matrix $\mathcal{A}$ kivela14; gomez2013. While the supra-adjacency matrix tells us which connections exist and how they are organized, the supra-Laplacian goes a step further; it captures how processes take place along those connections. It extends the analysis from merely identifying the links to understanding how different elements within the network influence each other. This transition from the static structure of connections to the dynamic interplay of influences is pivotal for understanding the overall behaviour of multiplex networks, setting the stage for a deeper exploration of the network’s properties through its eigenvalues sole2013.
@@ II.1.2 Structural Metrics
Once a multilayer network is defined, we need a toolkit of metrics to quantify its structure. While many metrics are generalizations from the ones in single-layer networks Aleta2020Nov, the most insightful ones are those that capture the unique features arising from the multi-layered structure (see Section II.1.4).
Starting with node-level metrics, the most basic measure of a node’s importance is its degree, $k$, defined as the number of connections it has. In multilayer networks, a node $i$ may have a different number of connections in each layer $\alpha$, yielding the per-layer degree $k_{i}^{\alpha}$. The multi-degree of a node is then the sum of its degrees across all layers. The degree is also a simple measure of the centrality of a node, in that it expresses its importance in the network. Over the last two decades, the list of centrality measures has considerably expanded, and each one of these metrics is defined according to some rules that limit their range of validity and interpretability. Multilayer networks are well-equipped with an extensive list of centrality measures, many of them being generalizations of the corresponding definitions in single-layer networks. Some of them are purely topological, such as the eigenvector, the Katz, and the betweenness centralities DeDomenico2013; DeDomenico2015ranking; sola2013eigenvector; taylor2021. Others are defined by means of dynamical processes. Notable examples include PageRank halu2013multiplex or a centrality related to the random walk occupation probability sole2016random, among others.
@@ II.1.3 Spectral properties
To characterize the network as a whole, spectral properties of the system’s matrix representations have been widely used. The eigenvalues of the Laplacian matrix of a network offer a wealth of information, which can be utilized to understand various dynamic properties of the network arenas2006; Almendral2007. The algebraic connectivity or Fiedler value Fiedler1973, given by the second-smallest eigenvalue ($\lambda_{2}$) of the supra-Laplacian matrix, is a key metric. A larger $\lambda_{2}$ indicates a more cohesive and robust network that is harder to break apart. Conversely, a small $\lambda_{2}$ can indicate potential points of vulnerability or fragmentation within the network. Higher eigenvalues $\{\lambda_{3},\lambda_{4},\,\ldots\}$ offer insights into more complex network structures and dynamicssole2013. For instance, they can be used to detect community structure within the network chauhan2009 (see Section II.2.1).
@@
Beyond structural insight, $\lambda_{2}$ plays a fundamental role in dynamical processes such as diffusion, consensus, and especially synchronization. In networks where each node represents a Kuramoto phase oscillator, $\lambda_{2}$ determines the timescale for convergence to the synchronized state Almendral2007. Small values of $\lambda_{2}$ are associated with bottlenecks or weakly connected components that can delay full synchronization. This spectral framework naturally extends to multiplex networks, where each node is present in multiple layers. Figure 1 shows the theoretical predictions for $\lambda_{2}$ of a two-layer network in the asymptotic regimes: $\lambda_{2}=2D_{x}$ for $D_{x}\ll 1$, and the superposition of both layers for $D_{x}\gg 1$, $D_{x}$ being the interlayer coupling. The plot also includes the eigenvalues of the individual layers and the corresponding supra-Laplacian eigenvalue.
@@
Figure 1: Comparison between the second smallest eigenvalues $\lambda_{2}$ of the different Laplacian matrices as a function of the interlayer coupling $D_{x}$. Image from gomez2013.
@@
Actually, through the behavior of $\lambda_{2}$ as a function of the interlayer connectivity, we can directly identify when multiplex networks will behave effectively as if each layer were an independent network, or instead when multiplexity kicks in with its emerging effects on the dynamics. This has been shown in a simplified model of a two-layered network where the interlayer connections are weighted by a homogeneous parameter $p$ in radicchi2013abrupt, but the arguments are general for networks with a larger number of layers. The algebraic connectivity of the supra-Laplacian $\mathcal{L}$ reads $\lambda_{2}(\mathcal{L})=2p$ for $p<p^{*}$ and is upper-bounded such that $\lambda_{2}(\mathcal{L})\leq\lambda_{2}(\mathcal{L_{A}}+\mathcal{L_{B}})/2$ for $p\geq p^{*}$, the critical point being $p^{*}\leq\lambda_{2}(\mathcal{L_{A}}+\mathcal{L_{B}})/4$. As it can be guessed graphically, but proved mathematically, the first derivative of $\lambda_{2}(\mathcal{L})$ is discontinuous. This emerges as a result of the crossing of two distinct groups of eigenvalues. Actually, these crossing events appear for the other eigenpairs of the graph Laplacian, with the exception of the smallest and largest eigenvalues. Consequently, this results in discontinuities in the first derivatives of the corresponding eigenvalues radicchi2013abrupt.
@@
The stability of the synchronized state in these systems can be analyzed using the Master Stability Function (MSF) formalism pecora98, which allows us to reduce the problem to a spectral one. The synchronizability of a network can be assessed by the eigenratio
@@
being $\lambda_{N}$ the largest eigenvalue of the Laplacian matrix. A smaller eigenratio $R$ implies greater synchronizability, since it indicates a wider range of coupling strengths for which the synchronized solution remains stable.
In sole2013, analytical approximations for the eigenratio $R$ were provided in two limiting regimes: weak interlayer coupling ($D_{x}\ll 1$) and strong interlayer coupling ($D_{x}\gg 1$). These approximations offer explicit insight into how the topology and interlayer coupling influence the dynamics. The predictions were shown to agree well with numerical results in a wide range of coupling values, as illustrated in Figure 2. The point at which the two analytical regimes intersect provides an estimate of the optimal coupling strength for synchronization performance.
Figure 2: Eigenratio $R=\lambda_{N}/\lambda_{2}$ as a function of the interlayer coupling strength $D_{x}$, adapted from sole2013. Circles represent exact numerical results, while the solid lines show the analytical approximations valid in the weak- and strong-coupling limits.

### v2

@@ Abstract
Multilayer network science has emerged as a central framework for analysing interconnected and interdependent complex systems. Its relevance has grown substantially with the increasing availability of rich, heterogeneous data, which makes it possible to uncover and exploit the inherently multilayered organisation of many real-world networks. In this review, we summarise recent developments in the field. On the theoretical and methodological front, we outline core concepts and survey advances in community detection, dynamical processes, temporal networks, higher-order interactions, and machine-learning-based approaches. On the application side, we discuss progress across diverse domains, including interdependent infrastructures, spreading dynamics, computational social science, economic and financial systems, ecological and climate networks, science-of-science studies, network medicine, and network neuroscience. We conclude with a forward-looking perspective, emphasizing the need for standardised datasets and software, deeper integration of temporal and higher-order structures, and a transition toward genuinely predictive models of complex systems.
@@ I Introduction
During the last 25 years, network science has established itself as a powerful and seamlessly integrated research discipline Dorogovtsev and Mendes (2013); Vespignani and Pastor-Satorras (2004); Caldarelli (2007); Barrat et al. (2008); Cohen and Havlin (2010); Newman (2018); Estrada (2011); Estrada and Knight (2015); Barabási (2016); Latora et al. (2017); Menczer et al. (2020); Bagrow and Ahn (2024). At variance with other, more niched disciplines, it has allowed us to enrich and to synergize with other branches of science: from sociology and economics to biology, chemistry, and physics. This has been possible because of the data revolution and increased computational power that took place in the twilight of the 20th century. Nowadays, applications of network science tools pervade all fields of science but, most importantly, the discipline has changed the way we think about interactions in natural and human-made complex systems.
@@
The momentum for the development of network science came with the discovery that seemingly diverse networks have universal properties that pervade many natural and artificial systems. Since 1998, ignited by the works of Watts and Strogatz and Barabási and Albert Watts and Strogatz (1998); Barabási and Albert (1999), systems as diverse as electric power grids, food chains, brain networks, protein networks, transcriptional networks, and social networks have been cast into the mathematical language of networks. We are now familiar with jargon that includes centrality, homogeneous and heterogeneous systems, communities, structural and dynamic phase transitions, weak and strong links, etc.
@@
Yet, as the field matured, it became increasingly clear that most real systems cannot be faithfully represented as individual networks. Many natural, social, and technological systems are instead composed of several types of interactions, giving rise to multilayer or interdependent structures Kivelä et al. (2014); Boccaletti et al. (2014); Battiston et al. (2014); Bianconi (2018); Aleta and Moreno (2019); Artime et al. (2022); De Domenico (2023). The conceptual roots of this idea trace back to the sociological notion of multiplexity, where individuals were known to maintain several types of social ties simultaneously within the same population Wasserman and Faust (1994). These early multi-relational studies anticipated many of the ideas later formalized in network science. Then, with the rise of large-scale digital data around 2010, empirical analyses began to demonstrate that distinct layers of interaction can exhibit markedly different structural and dynamical patterns, underscoring the need for explicit multilayer representations Szell et al. (2010).
@@
A major turning point came with the study of interdependent networks, where nodes in one layer depend on nodes in another layer for functionality Buldyrev et al. (2010); Vespignani (2010). This framework revealed that coupling multiple networks can fundamentally alter systemic behaviour, producing abrupt percolation transitions and cascading failures, phenomena not observed in isolated networks. Following these foundational insights, the field rapidly matured from focusing primarily on cascades and robustness to developing a comprehensive mathematical formalism and exploring a wide array of dynamic processes. Researchers worked to create a unified language and a robust mathematical framework to describe the diverse architectures of multilayer systems Kivelä et al. (2014); Boccaletti et al. (2014); Battiston et al. (2014), enabling the extension of classical concepts of network science to the multilayer context, such as centrality measures, among others. This theoretical consolidation sparked extensive research on dynamics across layers, including diffusion, synchronisation, epidemic spreading, and evolutionary games, revealing that interlayer coupling can either enhance or inhibit dynamical processes depending on structural correlations.
@@
In the following years, multilayer methodologies spread rapidly across disciplines. As we will discuss in the following sections, applications now span social, biological, and technological systems. Parallel advances in community detection and inference have made it possible to extract meaningful mesoscale organisation from high-dimensional, heterogeneous data. More recent developments integrate temporal, categorical, and hierarchical aspects within multilayer frameworks and connect them to higher-order network representations such as hypergraphs. Today, the multilayer paradigm stands as one of the most powerful generalizations in network science, bridging theoretical, computational, and empirical approaches to the study of interconnected complex systems.
@@ II.1.1 Definitions and Formalisms
A prominent example of a multilayer system is the coupling of a power grid and the internet. One layer represents the power grid, with nodes as power stations and substations, while links represent transmission lines, illustrating the distribution of electrical power. Meanwhile, the other layer represents the internet, with nodes as data centres, routing hubs, and servers, linked by data transmission pathways. This setup highlights the interdependency between layers: the internet needs power to function, while modern power grids rely on internet-based communication for their operation, something that can be encoded in the interlayer links. This multilayer network model is essential for studying cascading failures and the overall resilience of interconnected systems Buldyrev et al. (2010); De Domenico (2023).
A multiplex network is a specific and widely used type of multilayer network where nodes represent the same entity on every layer, and interlayer links only connect a node to itself across different layers. Each layer then represents a different type of relationship among the nodes. For instance, a social network can be modelled as a multiplex network where layers represent different platforms or social contexts Kivelä et al. (2014); Battiston et al. (2014).
To analyse multiplex networks, we need to generalize the adjacency matrix. Two primary formalisms are used. First, we have the tensor formalism, in which the separation of layers can be maintained explicitly. A multiplex network can then be represented by an adjacency tensor $\mathcal{M}$ of order 4 with elements $M_{ij}^{\alpha\beta}$. Here, $M_{ij}^{\alpha\alpha}$ represents the weight of the intralayer link between nodes $i$ and $j$ in layer $\alpha$, while $M_{ij}^{\alpha\beta}$ with $\alpha\neq\beta$ represents the interlayer connections. This is a more general and flexible representation, though it requires methods from tensor algebra for analysis De Domenico et al. (2013).
Second, we have the formalism based on the supra-adjacency matrix, which represents the entire multilayer system as a single, larger network. For a multiplex network with $N$ nodes and $L$ layers, the supra-adjacency matrix $\mathcal{A}$ is a $NL\times NL$ block matrix. The diagonal blocks are the adjacency matrices of each layer $\alpha$, $\text{{A}}^{\alpha}$, representing intralayer connections. The off-diagonal blocks encode the interlayer connections, which can be represented by the identity matrix I in the $\alpha\beta$ block if layer $\alpha$ and $\beta$ are connected, and 0 otherwise Kivelä et al. (2014). An example of supra-adjacency matrix of a multiplex network is
@@
Following these representations, we can also define the corresponding Laplacian matrices. For a single-layer graph, the Laplacian matrix is defined as $\text{{L}}=\text{{D}}-\text{{A}}$, where D is the diagonal matrix of node degrees. Similarly, the supra-Laplacian matrix $\mathcal{L}$ can be constructed from the supra-adjacency matrix $\mathcal{A}$ Kivelä et al. (2014); Gómez et al. (2013). While the supra-adjacency matrix tells us which connections exist and how they are organized, the supra-Laplacian goes a step further; it captures how processes take place along those connections. It extends the analysis from merely identifying the links to understanding how different elements within the network influence each other. This transition from the static structure of connections to the dynamic interplay of influences is pivotal for understanding the overall behaviour of multiplex networks, setting the stage for a deeper exploration of the network’s properties through its eigenvalues Solé-Ribalta et al. (2013).
@@ II.1.2 Structural Metrics
Once a multilayer network is defined, we need a toolkit of metrics to quantify its structure. While many metrics are generalizations from the ones in single-layer networks Aleta et al. (2020b), the most insightful ones are those that capture the unique features arising from the multi-layered structure (see Section II.1.4).
Starting with node-level metrics, the most basic measure of a node’s importance is its degree, $k$, defined as the number of connections it has. In multilayer networks, a node $i$ may have a different number of connections in each layer $\alpha$, yielding the per-layer degree $k_{i}^{\alpha}$. The multi-degree of a node is then the sum of its degrees across all layers. The degree is also a simple measure of the centrality of a node, in that it expresses its importance in the network. Over the last two decades, the list of centrality measures has considerably expanded, and each one of these metrics is defined according to some rules that limit their range of validity and interpretability. Multilayer networks are well-equipped with an extensive list of centrality measures, many of them being generalizations of the corresponding definitions in single-layer networks. Some of them are purely topological, such as the eigenvector, the Katz, and the betweenness centralities De Domenico et al. (2013, 2015b); Solá et al. (2013); Taylor et al. (2021). Others are defined by means of dynamical processes. Notable examples include PageRank Halu et al. (2013) or a centrality related to the random walk occupation probability Solé-Ribalta et al. (2016), among others.
@@ II.1.3 Spectral properties
To characterise the network as a whole, spectral properties of the system’s matrix representations have been widely used. The eigenvalues of the Laplacian matrix of a network offer a wealth of information, which can be utilized to understand various dynamic properties of the network Arenas et al. (2006); Almendral and Díaz-Guilera (2007). The algebraic connectivity or Fiedler value Fiedler (1973), given by the second-smallest eigenvalue ($\lambda_{2}$) of the supra-Laplacian matrix, is a key metric. A larger $\lambda_{2}$ indicates a more cohesive and robust network that is harder to break apart. Conversely, a small $\lambda_{2}$ can indicate potential points of vulnerability or fragmentation within the network. Higher eigenvalues $\{\lambda_{3},\lambda_{4},\,\ldots\}$ offer insights into more complex network structures and dynamicsSolé-Ribalta et al. (2013). For instance, they can be used to detect community structure within the network Chauhan et al. (2009) (see Section II.2.1).
@@
Beyond structural insight, $\lambda_{2}$ plays a fundamental role in dynamical processes such as diffusion, consensus, and especially synchronisation. In networks where each node represents a Kuramoto phase oscillator, $\lambda_{2}$ determines the timescale for convergence to the synchronised state Almendral and Díaz-Guilera (2007). Small values of $\lambda_{2}$ are associated with bottlenecks or weakly connected components that can delay full synchronisation. This spectral framework naturally extends to multiplex networks, where each node is present in multiple layers. Figure 1 shows the theoretical predictions for $\lambda_{2}$ of a two-layer network in the asymptotic regimes: $\lambda_{2}=2D_{x}$ for $D_{x}\ll 1$, and the superposition of both layers for $D_{x}\gg 1$, $D_{x}$ being the interlayer coupling. The plot also includes the eigenvalues of the individual layers and the corresponding supra-Laplacian eigenvalue.
@@
Figure 1: Comparison between the second smallest eigenvalues $\lambda_{2}$ of the different Laplacian matrices as a function of the interlayer coupling $D_{x}$. Reprinted figure with permission from Gómez et al. (2013). ©2013 by the American Physical Society.
@@
Actually, through the behaviour of $\lambda_{2}$ as a function of the interlayer connectivity, we can directly identify when multiplex networks will behave effectively as if each layer were an independent network, or instead when multiplexity kicks in with its emerging effects on the dynamics. This has been shown in a simplified model of a two-layered network where the interlayer connections are weighted by a homogeneous parameter $p$ in Radicchi and Arenas (2013), but the arguments are general for networks with a larger number of layers. The algebraic connectivity of the supra-Laplacian $\mathcal{L}$ reads $\lambda_{2}(\mathcal{L})=2p$ for $p<p^{*}$ and is upper-bounded such that $\lambda_{2}(\mathcal{L})\leq\lambda_{2}(\mathcal{L_{A}}+\mathcal{L_{B}})/2$ for $p\geq p^{*}$, the critical point being $p^{*}\leq\lambda_{2}(\mathcal{L_{A}}+\mathcal{L_{B}})/4$. As it can be guessed graphically, but proved mathematically, the first derivative of $\lambda_{2}(\mathcal{L})$ is discontinuous. This emerges as a result of the crossing of two distinct groups of eigenvalues. Actually, these crossing events appear for the other eigenpairs of the graph Laplacian, with the exception of the smallest and largest eigenvalues. Consequently, this results in discontinuities in the first derivatives of the corresponding eigenvalues Radicchi and Arenas (2013).
@@
The stability of the synchronised state in these systems can be analysed using the Master Stability Function (MSF) formalism Pecora and Carroll (1998), which allows us to reduce the problem to a spectral one. The synchronisability of a network can be assessed by the eigenratio
@@
where $\lambda_{N}$ is the largest eigenvalue of the Laplacian matrix. A smaller eigenratio $R$ implies greater synchronisability, since it indicates a wider range of coupling strengths for which the synchronised solution remains stable.
In Solé-Ribalta et al. (2013), analytical approximations for the eigenratio $R$ were provided in two limiting regimes: weak interlayer coupling ($D_{x}\ll 1$) and strong interlayer coupling ($D_{x}\gg 1$). These approximations offer explicit insight into how the topology and interlayer coupling influence the dynamics. The predictions were shown to agree well with numerical results in a wide range of coupling values, as illustrated in Figure 2. The point at which the two analytical regimes intersect provides an estimate of the optimal coupling strength for synchronisation performance.
Figure 2: Eigenratio $R=\lambda_{N}/\lambda_{2}$ as a function of the interlayer coupling strength $D_{x}$. Circles represent exact numerical results, while the solid lines show the analytical approximations valid in the weak- and strong-coupling limits. Adapted figure with permission from Solé-Ribalta et al. (2013). ©2013 by the American Physical Society.